# Q-Bench-Video: Benchmarking the Video Quality Understanding of LMMs

### v1

@@ Abstract
With the rising interest in research on Large Multi-modal Models (LMMs) for video understanding, many studies have emphasized general video comprehension capabilities, neglecting the systematic exploration into video quality understanding. To address this oversight, we introduce Q-Bench-Video in this paper, a new benchmark specifically designed to evaluate LMMs’ proficiency in discerning video quality. a) To ensure video source diversity, Q-Bench-Video encompasses videos from natural scenes, AI-generated Content (AIGC), and Computer Graphics (CG). b) Building on the traditional multiple-choice questions format with the Yes-or-No and What-How categories, we include Open-ended questions to better evaluate complex scenarios. Additionally, we incorporate the video pair quality comparison question to enhance comprehensiveness. c) Beyond the traditional Technical, Aesthetic, and Temporal distortions, we have expanded our evaluation aspects to include the dimension of AIGC distortions, which addresses the increasing demand for video generation. Finally, we collect a total of 2,378 question-answer pairs and test them on 12 open-source & 5 proprietary LMMs. Our findings indicate that while LMMs have a foundational understanding of video quality, their performance remains incomplete and imprecise, with a notable discrepancy compared to human performance. Through Q-Bench-Video, we seek to catalyze community interest, stimulate further research, and unlock the untapped potential of LMMs to close the gap in video quality understanding.
@@ 1 Introduction
As the field of artificial intelligence (AI) continues to evolve, Large Multi-modal Models (LMMs) (Ye et al., 2024; Li et al., 2024a; Chen et al., 2024; Ke et al., 2023; Xu et al., 2024) are progressively utilized in high-level video understanding tasks. These models have shown remarkable capabilities in analyzing and interpreting the semantic content of videos, such as classifying objects, identifying actions, and recognizing events. However, the aspect of video quality, which is vital for optimizing compression and transmission systems, enhancing viewer experience, and establishing standards for high-quality video generation, has received less attention. Although numerous LMM video benchmarks (Fu et al., 2024; Fang et al., 2024; Wu et al., 2024a) have been developed to assess the semantic understanding of videos by LMMs comprehensively, benchmarks systematically targeting video quality are still lacking. Additionally, while semantic understanding is closely linked to high-level video information, the perception and understanding of low-level information are crucial in video quality (Chikkerur et al., 2011; Li et al., 2024b). Thus, current video benchmarks fail to adequately evaluate the video quality understanding capabilities of LMMs.
To address this gap, we introduce Q-Bench-Video, a novel benchmark specifically designed to systematically evaluate the video quality understanding of LMMs. As illustrated in Fig. 1, our benchmark encompasses a wide range of video content, including natural scenes, AI-generated Content (AIGC), and Computer Graphics (CG), ensuring diversity in video sources. In addition, to maintain a reasonable distribution of source video quality, we employ uniform sampling from video datasets that contain subjective quality annotations. This approach guarantees comprehensive coverage of the quality spectrum while avoiding imbalanced quality distributions. Moreover, we extend beyond traditional video evaluations by incorporating both multiple-choice questions (MCQs) and open-ended questions. This enables a more thorough analysis of LMMs’ ability to discern video quality across diverse scenarios. We further introduce a new evaluation dimension specifically tailored to assess distortions related to AIGC, which are increasingly prominent in video generation tasks. Recognizing the importance of quality comparison settings in real-world applications, such as camera parameter optimization and AIGC video generation, we further incorporate video pairs to facilitate quality comparison assessment as well. In total, we collect 1800 videos and annotated 2,378 question-answer pairs for validation, creating a robust framework for systematic evaluation.
Through the rigorous experiment, we demonstrate that while LMMs show promise in video quality assessment, their performance lags significantly behind human-level understanding. By offering a systematic and thorough evaluation of LMMs’ video quality perception, Q-Bench-Video aims to foster research in this underexplored area and push the boundaries of LMM capabilities.
@@
Our contributions can be summarized as follows:
  • We introduce Q-Bench-Video, the first comprehensive benchmark explicitly designed to assess the video quality understanding capabilities of LMMs. This benchmark includes a diverse collection of source videos and ensures a balanced quality distribution, complemented by human-crafted question-answer annotations to enable thorough evaluation.
@@
Table 1: Overview of the diverse video source datasets in the Q-Bench-Video. We consider various video content types, including natural scenes, AIGC, and CG videos. The term ‘MOS’ denotes that the videos are annotated via Mean Opinion Scores under ITU standards (itu, 2000). We have conducted uniform sampling based on their quality labels to ensure a balanced quality distribution.
@@
Natural (1000) LSVQ (Ying et al., 2021) Spatial & Temporal 600 39K
MaxWell (Wu et al., 2023b) Spatial & Temporal & Aesthetic 350 4.5K
WaterlooSQoE-III (Duanmu et al., 2018) Quality-of-Experience 20 450
WaterlooSQoE-IV (Duanmu et al., 2020) Quality-of-Experience 30 1,350
@@
VideoFeedback (He et al., 2024b) $\times$ Quality & Text Alignment 400 37.6K
@@ 2 Related Works
2.1 Large Multi-modal Video Models and Benchmarks
The rapid advancement of Large Multi-modal Models (LMMs) in recent years (Liu et al., 2023b; a; 2024a; Chen et al., 2024; Zhang et al., 2023a; Ye et al., 2023a) has showcased their remarkable perception and cognitive abilities across various multimodal benchmarks for images (Liu et al., 2023d; Wu et al., 2024b; Fu et al., 2023; Zhang et al., 2024d; Marino et al., 2019). As development progresses, the focus of visual analysis has gradually shifted from images to videos. Early efforts (Li et al., 2023a; Lin et al., 2023; Liu et al., 2023c; Xu et al., 2024) aimed at unlocking the video understanding potential of LMMs have yielded promising results. However, initial video-based benchmarks (Li et al., 2023b; Wang et al., 2023; Mangalam et al., 2023) typically concentrated on specific aspects of video comprehension, falling short of fully capturing the performance of these models due to limitations such as a lack of diversity in video types and inadequate coverage of temporal dynamics. In response, more recent video benchmarks (Fu et al., 2024; Fang et al., 2024) have moved toward a more comprehensive evaluation of LMMs. Nonetheless, these efforts primarily focus on high-level semantic understanding without systematic exploration of video quality.
@@ 2.2 Video Quality Assessment
Video Quality Assessment (VQA) is a task aimed at quantifying video scores based on visual quality. Initially, early VQA methods employ hand-crafted features extracted from videos and regress the features into quality scores (Zheng et al., 2022; Vu et al., 2011; Li et al., 2018). With the emergence of deep neural networks, a shift occurr as numerous methods adopted deep learning techniques for VQA tasks (Li et al., 2019; Sun et al., 2022; Li et al., 2022; Wen et al., 2024). As the field progresses, newer methods begin incorporating considerations for both temporal dynamics and aesthetic qualities, leading to a more holistic approach to video quality analysis (Wu et al., 2022a; 2023a; Zhang et al., 2023b; Ahn & Lee, 2018; He et al., 2024a). Moreover, the evolution of large-scale models has further revolutionized VQA methodologies. Many recent approaches have redefined the traditional quality assessment process into a quality question-answering format (Wu et al., 2024c; Ge et al., 2024; Zhang et al., 2024b). This adaptation leverages the substantial prior knowledge embedded in large models to enhance the precision of quality quantification (Zhang et al., 2024c). Despite these technological advances, VQA still grapples with challenges in providing interpretable quality scores and deepening the understanding of how models perceive and analyze video quality.
@@ 3.1 Benchmark Principle
The Q-Bench-Video is designed based on three guiding principles: (1) It encompasses a broad spectrum of video content, including natural scenes, AIGC, and CG videos. (2) It ensures a comprehensive and representative sampling process across a wide quality range, enhancing the benchmark’s overall effectiveness. (3) It primarily focuses on the aspects of video quality that significantly influence the viewing experience, including technical, aesthetic, temporal, and AIGC distortions. This significantly differs from other video benchmarks that prioritize semantic understanding. Additionally, the video pair quality comparison is integrated to address the challenges associated with comparing video quality. The construction can be overviewed in Fig. 1.
@@ 3.2 Source Videos Collection
As shown in Table 1, the source videos are primarily gathered from video quality assessment datasets. We selected videos from these datasets for two main reasons: (1) These datasets have inherently considered the diversity of quality features during video selection; (2) These datasets possess quality annotations that adhere to ITU standards (itu, 2000) (with the exception of VideoFeedback), which allow us to accurately and authentically sample videos.
Our sampling method primarily employs a uniform approach, extracting videos evenly from each dataset based on the quality range. Moreover, considering the current popularity of AIGC and CG videos, we have also incorporated a selection of these video types. For a detailed description of the datasets and the sampling procedure, please refer to Appendix A and Appendix B.
@@ 3.3.1 Question Types
Yes-or-No Questions. The basic Yes-or-No questions are designed to prompt LMMs to make binary judgments on video quality queries, typically limited to the answers Yes or No. To address the potential bias in LMMs that may skew towards yes responses, we employ a rigorous annotation process. This process ensures that the distribution of correct answers, either Yes or No, remains balanced at about 50%/50% ratio (see Appendix D). This balanced approach allows for a more accurate assessment of LMMs’ performance on Yes-or-No questions.
What-How Questions. The What-How questions are commonly utilized in benchmarks for LMMs. The What questions focus on identifying specific distortions (e.g., What is the most apparent distortion in this video?). On the other hand, the How questions are employed to distinguish the finer details of distortion levels (e.g., How is the overall clarity of this video?). Including both What and How questions allows Q-Bench-Video to thoroughly and meticulously evaluate LMMs’ ability on identifying video distortions and evaluating the distortion levels.
Open-ended Questions. It’s important to note that the two types of questions previously mentioned require LMMs to select the correct answer from a predefined set. However, in many real-world scenarios, Open-ended Questions, which do not restrict responses to a predefined set, are often more necessary and challenging for LMMs (e.g., What are the possible factors that lead to the low clarity of this video? Please list and explain.). By adopting this form of questioning, we can better assess an LMM’s ability to perceive video quality in real-world conditions.
@@ 3.3.2 Quality Concerns
Technical Distortions. Technical distortions refer to the low-level degradation in video quality that arises from the limitations of recording, compression, and transmission (Su et al., 2021; Ying et al., 2021). These distortions often include artifacts such as blurring, noise, compression artifact, exposure, etc., which are directly tied to the technical processes used in video production and delivery.
Aesthetic Distortions. Aesthetic distortions involve deviations from the intended visual style, artistic design, or creative intent that negatively affect the viewer’s perception of the video (Wu et al., 2022b; Huang et al., 2024). These distortions can include aspects such as confusing color, poor composition, lighting inconsistencies, or distracting elements that reduce the overall aesthetic appeal. Unlike technical distortions, aesthetic distortions are subjective and might be affected by viewer preference, cultural context, or artistic norms.
Temporal Distortions. Temporal distortions are related to the degradation of visual quality over time, impacting the fluidity and consistency of the video (Seshadrinathan & Bovik, 2009). These distortions manifest as issues like screen shake, flickering, motion inconsistency, frame drops, and stuttering that result from unstable shooting devices, dramatically changing lighting conditions, and unstable bitrate environments (Wu et al., 2023b). Such disruptions hinder the viewer’s natural perception of the video, leading to a disjointed and unpleasant viewing experience.
AIGC Distortions. AIGC distortions pertain to imperfections and unnaturalness specifically arising from the generation of video content through AI models (Liu et al., 2024b; Zhang et al., 2024c). These distortions may include unnatural textures, inconsistent lighting, uncanny facial features, or unrealistic object behavior that result from limitations or biases in the training data, model architecture, or generative process. These distortions are unique to AI-generated content and require specialized evaluation metrics that consider both the technical and perceptual quality aspects.
@@ 3.3.3 Single Videos & Video Pairs
Accurately comparing and jointly analyzing the quality of video pairs is sometimes more crucial than assessing the quality of a single video, especially in scenarios such as performance tests in video compression and quality control in video generation (In which it is more important to find out Which video is better in visual quality and why?(Zhu et al., 2024; Wu et al., 2024d). Therefore, in Q-Bench-Video, we include video quality queries for both single videos and video pairs.
Single Videos. Queries related to single videos can primarily be categorized into two types: a) Global perception, which involves questions about the overall visual quality of the video, such as How is the overall contrast of this video? b) Referring perception, which focuses on the visual quality of specific elements within the video, like querying What is the most apparent distortion when the player strikes the ball? Through these approaches, we aim to comprehensively evaluate the LMMs’ ability to perceive both the overall and localized aspects of video quality.
Video Pairs. Firstly, to ensure that the comparison of video pairs is clear and meaningful, comparisons are only made between videos from the same source, such as videos from natural sources being paired together while CG videos and AIGC videos are not being paired. There are mainly two types of video pair categories: a) Joint analysis, which involves understanding the shared quality features of the video pairs, for example, such as asking Are both videos blurry? b) Comparative analysis, which involves comparing the quality dimension across two videos, such as How does the level of brightness in the first video compare to that in the second one? It is important to note that we further categorize comparisons based on the difference in quality labels of the videos involved into coarse-grain (with relatively more significant visual quality differences) and fine-grain (with relatively minor visual quality differences) comparisons. More details can be found in Appendix J.

### v2

# Q-Bench-Video: Benchmark the Video Quality Understanding of LMMs

@@ Abstract
With the rising interest in research on Large Multi-modal Models (LMMs) for video understanding, many studies have emphasized general video comprehension capabilities, neglecting the systematic exploration into video quality understanding. To address this oversight, we introduce Q-Bench-Video in this paper, a new benchmark specifically designed to evaluate LMMs’ proficiency in discerning video quality. a) To ensure video source diversity, Q-Bench-Video encompasses videos from natural scenes, AI-generated content (AIGC), and computer graphics (CG). b) Building on the traditional multiple-choice questions format with the Yes-or-No and What-How categories, we include Open-ended questions to better evaluate complex scenarios. Additionally, we incorporate the video pair quality comparison question to enhance comprehensiveness. c) Beyond the traditional Technical, Aesthetic, and Temporal distortions, we have expanded our evaluation aspects to include the dimension of AIGC distortions, which addresses the increasing demand for video generation. Finally, we collect a total of 2,378 question-answer pairs and test them on 12 open-source & 5 proprietary LMMs. Our findings indicate that while LMMs have a foundational understanding of perceptual video quality, their performance remains incomplete and imprecise, with a notable discrepancy compared to the performance of human beings. Through Q-Bench-Video, we seek to catalyze community interest, stimulate further research, and unlock the untapped potential of LMMs to close the gap in video quality understanding.
@@ 1 Introduction
As the field of artificial intelligence (AI) continues to evolve, Large Multi-modal Models (LMMs) [52, 19, 4, 16, 50] are progressively utilized in high-level video understanding tasks. These models have shown remarkable capabilities in analyzing and interpreting the semantic content of videos, such as classifying objects, identifying actions, and recognizing events. However, the aspect of video quality, which is vital for optimizing compression and transmission systems, enhancing viewer experience, and establishing standards for high-quality video generation, has received less attention. Although numerous LMM video benchmarks[10, 8, 23] have been developed to assess the semantic understanding of videos by LMMs comprehensively, benchmarks systematically targeting video quality are still lacking. Additionally, while semantic understanding is closely linked to high-level video information, the perception and understanding of low-level information are crucial in video quality [5, 25] as well.
@@
To address this gap, we introduce Q-Bench-Video, a novel benchmark specifically designed to systematically evaluate the video quality understanding of LMMs. As illustrated in Fig. 1, our benchmark encompasses a wide range of video content, including natural scenes, AI-generated content (AIGC), and computer graphics (CG), ensuring diversity in video sources. In addition, to maintain a reasonable distribution of source video quality, we employ uniform sampling from video datasets that contain subjective quality annotations. This approach guarantees comprehensive coverage of the quality spectrum while avoiding imbalanced quality distributions. Moreover, we extend beyond traditional video evaluations by incorporating both multiple-choice questions (MCQs) and open-ended questions. This enables a more thorough analysis of LMMs’ ability to discern video quality across diverse scenarios. We further introduce a new evaluation dimension specifically tailored to assess distortions related to AIGC, which are increasingly prominent in video generation tasks. Recognizing the importance of quality comparison settings in real-world applications, such as camera parameter optimization and AIGC video generation, we further incorporate video pairs to facilitate quality comparison assessment as well. In total, we collect 1800 videos and annotated 2,378 question-answer pairs for validation, creating a robust framework for systematic evaluation on video quality understanding.
Through the rigorous experiment, we demonstrate that while LMMs show promise in video quality assessment, their performance lags significantly behind human-level understanding. By offering a systematic and thorough evaluation of LMMs’ video quality perception, Q-Bench-Video aims to foster research in this underexplored area and push the boundaries of LMM capabilities on video quality understanding. Our contributions can be summarized as follows:
  • We introduce Q-Bench-Video, the first comprehensive benchmark explicitly designed to assess the video quality understanding capabilities of LMMs. This benchmark includes a diverse collection of source videos and ensures a balanced quality distribution, complemented by human-crafted question-answer annotations.
@@
Table 1: Overview of the diverse video source datasets in the Q-Bench-Video. We consider various video content types, including natural scenes, AIGC, and CG videos. The term ‘MOS’ denotes that the videos are annotated via Mean Opinion Scores under ITU standards [1]. We have conducted uniform sampling based on their quality labels to ensure a balanced quality distribution.
@@
      Natural (1000)       LSVQ [55]       ✓       Spatial & Temporal       600       39K
      MaxWell [46]       ✓       Spatial & Temporal & Aesthetic       350       4.5K
      WaterlooSQoE-III [6]       ✓       Quality-of-Experience       20       450
      WaterlooSQoE-IV [7]       ✓       Quality-of-Experience       30       1,350
@@
      VideoFeedback [13]       $\times$       Quality & Text Alignment       400       37.6K
@@ 2 Related Works
2.1 Video LMMs and Benchmarks
The rapid advancement of Large Multi-modal Models (LMMs) in recent years [29, 28, 30, 4, 57, 53, 68] has showcased their remarkable perception and cognitive abilities across various multimodal benchmarks for images [33, 47, 9, 63, 35]. As development progresses, the focus of visual analysis has gradually shifted from images to videos. Early efforts [22, 27, 31, 50] aimed at unlocking the video understanding potential of LMMs have yielded promising results. However, initial video-based benchmarks [24, 41, 34] typically concentrated on specific aspects of video comprehension, falling short of fully capturing the performance of these models due to limitations such as a lack of diversity in video types and inadequate coverage of temporal dynamics. In response, recent video benchmarks [10, 8] have moved toward a more comprehensive evaluation of LMMs. Nonetheless, these efforts focus on high-level semantic understanding without systematic exploration of video quality.
@@ 2.2 Video Quality Assessment
Video Quality Assessment (VQA) is a task aimed at quantifying video scores based on visual quality. Initially, early VQA methods employ hand-crafted features extracted from videos and regress the features into quality scores [65, 40, 26, 59, 17, 66, 20]. With the emergence of deep neural networks, a shift occurs as numerous methods adopted deep learning techniques for VQA tasks [21, 38, 18, 42]. As the field progresses, newer methods begin incorporating considerations for both temporal dynamics and aesthetic qualities, leading to a more holistic approach to video quality analysis [43, 45, 60, 3, 12]. Moreover, the evolution of large-scale models has further revolutionized VQA methodologies. Many recent approaches have redefined the traditional quality assessment process into a quality question-answering format [48, 11, 61, 15, 64]. This adaptation leverages the substantial prior knowledge embedded in large models to enhance the precision of quality quantification [62]. Despite these technological advances, VQA still grapples with challenges in providing interpretable quality scores and deepening the understanding of how models perceive and analyze video quality more comprehensively.
@@ 3.1 Benchmark Principle
The Q-Bench-Video is designed based on three guiding principles: (1) It encompasses a broad spectrum of video content, including natural scenes, AIGC, and CG videos. (2) It ensures a comprehensive and representative sampling process across a wide quality range, enhancing the benchmark’s overall effectiveness. (3) It primarily focuses on the aspects of video quality that significantly influence the viewing experience, including technical, aesthetic, temporal, and AIGC distortions. This significantly differs from other video benchmarks that prioritize semantic understanding. Additionally, the video pair quality comparison is integrated to address the challenges associated with comparing video quality. The overall construction process can be clearly overviewed in Fig. 1.
@@ 3.2 Source Videos Collection
As shown in Table 1, the source videos are primarily gathered from video quality assessment datasets. We selected videos from these datasets for two main reasons: (1) These datasets have inherently considered the diversity of quality features during video selection; (2) These datasets possess quality annotations that adhere to ITU standards [1] (with the exception of VideoFeedback), which allow us to accurately and authentically sample videos.
Our sampling method primarily employs a uniform approach, extracting videos evenly from each dataset based on the quality range. Moreover, considering the current popularity of AIGC and CG videos, we have also incorporated a selection of these video types. (More details in the Supp.)
@@ 3.3.1 Question Types
I) Yes-or-No Questions. The basic Yes-or-No questions are designed to prompt LMMs to make binary judgments on video quality queries, typically limited to the answers Yes or No. To address the potential bias in LMMs that may skew towards yes responses, we employ an annotation adjustment process, which modifies the questions and correct answers if necessary (based on the Yes-or-No ratio). This process ensures that the distribution of correct answers, either Yes or No, remains balanced at about 50%:50% ratio in the end. This balanced approach allows for a more accurate assessment of LMMs’ performance on Yes-or-No questions.
II) What-How Questions. The What-How questions are commonly utilized in benchmarks for LMMs. The What questions focus on identifying specific distortions (e.g., What is the most apparent distortion in this video?). On the other hand, the How questions are employed to distinguish the finer details of distortion levels (e.g., How is the overall clarity of this video?). Including both What and How questions allows Q-Bench-Video to thoroughly and meticulously evaluate LMMs’ ability on identifying video distortions and evaluating the distortion levels.
III) Open-ended Questions. It’s important to note that the two types of questions previously mentioned require LMMs to select the correct answer from a predefined set. However, in many real-world scenarios, Open-ended Questions, which do not restrict responses to a predefined set, are often more necessary and challenging for LMMs (e.g., What are the possible factors that lead to the low clarity of this video? Please list and explain.). By adopting this form of questioning, we can better assess an LMM’s ability to perceive video quality in real-world conditions.
@@ 3.3.2 Quality Concerns
I) Technical Distortions. Technical distortions refer to the low-level degradation in video quality that arises from the limitations of recording, compression, and transmission [37, 55]. These distortions often include artifacts such as blurring, noise, compression artifact, exposure, etc., which are directly tied to the technical processes used in video production and delivery. The CG-specific distortions related to geometry and texture are taken into consideration.
II) Aesthetic Distortions. Aesthetic distortions involve deviations from the intended visual style, artistic design, or creative intent that negatively affect the viewer’s perception of the video [44, 14]. These distortions can include aspects such as confusing color, poor composition, lighting inconsistencies, or distracting elements that reduce the overall aesthetic appeal. Unlike technical distortions, aesthetic distortions are subjective and might be affected by viewer preference, cultural context, or artistic norms.
III) Temporal Distortions. Temporal distortions are related to the degradation of visual quality over time, impacting the fluidity and consistency of the video [36]. These distortions manifest as issues like screen shake, flickering, motion inconsistency, frame drops, and stuttering that result from unstable shooting devices, dramatically changing lighting conditions, and unstable bitrate environments [46]. Such disruptions hinder the viewer’s natural perception of the video, leading to a disjointed and unpleasant viewing experience, especially in live-streaming scenarios.
IV) AIGC Distortions. AIGC distortions pertain to imperfections and unnaturalness specifically arising from the generation of video content through AI models [32, 62]. These distortions may include unnatural textures, inconsistent lighting, uncanny facial features, or unrealistic object behavior that result from limitations or biases in the training data, model architecture, or generative process. These distortions are unique to AI-generated content and require specialized evaluation metrics that consider both the technical and perceptual quality aspects.
((a))
((b))
@@ 3.3.3 Single Videos & Video Pairs
Accurately comparing and jointly analyzing the quality of video pairs is sometimes more crucial than assessing the quality of a single video, especially in scenarios such as performance tests in video compression and quality control in video generation (In which it is more important to find out Which video is better in visual quality and why?[67, 49]. Therefore, in Q-Bench-Video, we include video quality queries for both single videos and video pairs.
((a))
((b))
I) Single Videos. Queries related to single videos can primarily be categorized into two types: a) Global perception, which involves questions about the overall visual quality of the video, such as How is the overall contrast of this video? b) Referring perception, which focuses on the visual quality of specific elements within the video, like querying What is the most apparent distortion when the player strikes the ball? Through these approaches, we aim to comprehensively evaluate the LMMs’ ability to perceive both the overall and localized aspects of video quality.
II) Video Pairs. Firstly, to ensure that the comparison of video pairs is clear and meaningful, comparisons are only made between videos from the same source, such as videos from natural sources being paired together while CG videos and AIGC videos are not being paired. There are mainly two types of video pair categories: a) Joint analysis, which involves understanding the shared quality features of the video pairs, for example, such as asking Are both videos blurry? b) Comparative analysis, which involves comparing the quality dimension across two videos, such as How does the level of brightness in the first video compare to that in the second one? It is important to note that we further categorize comparisons based on the difference in quality labels of the videos involved into coarse-grain (with relatively more significant visual quality differences) and fine-grain (with relatively minor visual quality differences) comparisons. (More details in the Supp.)